Find a general solution for the differential equation -y''=6x+xe^x (a) using the method of undetermined coefficients. (b) using variation of parameters

1
Expert's answer
2022-06-02T13:50:40-0400

a)

Homogeneous differential equation is


"y''=0"

Characteristic (auxiliary) equation


"r^2=0"

The general solution of the homogeneous differential equation is


"y_h=C_1+C_2x"


Find the particular solution of the non homogeneous differential equation in the form


"y_p=x^2(Ax+B)+(Cx+D)e^x"

Then


"y_p'=3Ax^2+2Bx+Ce^x+(Cx+D)e^x"

"y_p''=6Ax+2B+2Ce^x+Cxe^x+De^x"

Substitute


"6Ax+2B+2Ce^x+Cxe^x+De^x=-6x-xe^x"

"A=-1, B=0, C=-1, D=2"

"C=-1"

The general solution of the given nonhomogeneous differential equation is

"y=-x^3-xe^x+2e^x+C_1+C_2x"

b)

Homogeneous differential equation is


"y''=0"

Characteristic (auxiliary) equation


"r^2=0"

The general solution of the homogeneous differential equation is


"y_h=C_1+C_2x"

"y'=C_1'+C_2'x+C_2"

Let

"C_1'+C_2'x=0"

Then


"y'=C_2"

"y''=C_2'"

Substitute


"C_2'=-6x-xe^x"

"C_2=\int(-6x-xe^x)dx"

"\int xe^xdx=xe^x-\int e^xdx=xe^x-e^x-C_3"

"C_2=-3x^2-xe^x+e^x+C_3"


"C_1'=-C_2'x"

"C_1'=6x^2+x^2e^x"

"C_1=\int(6x^2+x^2e^x)dx"

"\int x^2e^xdx=x^2e^x-2\int xe^xdx"

"=x^2e^x-2xe^x+2e^x+C_4"


"C_1=2x^3+x^2e^x-2xe^x+2e^x+C_4"

The general solution of the given nonhomogeneous differential equation is


"y=2x^3+x^2e^x-2xe^x+2e^x+C_4"

"-3x^3-x^2e^x+xe^x+C_3x"

"y=-x^3-xe^x+2e^x+C_4+C_3x"